# FFT-based Dynamic Subspace Selection for Low-Rank Adaptive Optimization of Large Language Models

### v2

@@ Abstract
Low-rank optimization has emerged as a promising direction in training large language models (LLMs) to improve running time and reduce the memory usage of adaptive optimizers by constraining learning to a lower-dimensional space. Prior work typically projects gradients of linear layers using approaches based on Singular Value Decomposition (SVD) or QR-decomposition. Applying these techniques individually to each layer in large models is computationally expensive and incurs additional memory costs due to storing the projection matrices. In this work, we propose a computationally efficient and conceptually simple, two-step procedure to approximate SVD/QR-based gradient projections into lower-dimensional spaces by using a predefined orthogonal matrix of the Discrete Cosine Transform (DCT). We dynamically select columns from the DCT matrix based on their alignment with the gradient of each layer. The effective projection matrices are obtained via a simple matmul with the DCT matrix in $O(n^{3})$ time, followed by a lightweight sorting step to identify the most relevant basis vectors. For large layers, DCT can be computed via Makhoul’s $N$-point algorithm based on Fast Fourier Transform (FFT) in $O(n^{2}\log(n))$ time. Due to the predefined nature of the orthogonal bases, they are computed once at the start of training. Our numerical experiments on both pre-training and fine-tuning tasks demonstrate the effectiveness of our dual strategy in approximating optimal low-rank projections, obtaining an approach with rank-independent runtime that matches the performance of costly SVD/QR-based methods while achieving faster runtime and reduced memory usage by up to $25\%$ across different model sizes.
@@ 1 Introduction
Recently, the Muon (Jordan et al., 2024) optimizer sparked the community’s attention due to the faster convergence in pretraining settings, achieved by orthogonalizing the momentum matrix using Newton-Schulz Newton-Schulz, an iterative procedure difficult to parallelize for large-scale pretraining runs as the full-sized matrices must be materialized on GPU. The Dion optimizer (Ahn et al., 2025) aims to reduce this overhead by employing low-rank, orthogonal updates. However, it requires storing a projection matrix for each layer and uses QR-decomposition to orthogonalize the low-rank components, making its running time dependent on the rank.
@@
  • We motivate that DCT is a good candidate for our dynamic column selection approach due to the reduced time complexity to compute the alignments with the input matrix via the Makhoul’s $N$-point algorithm to compute a fast DCT with $O(n^{2}\log(n))$ time complexity compared to $O(n^{3})$ incurred by a basic matrix multiplication (Makhoul, 1980). This can yield speedups of up to $8-50\times$ in computing the alignments for large layers (Appendix C).
@@
  • We show the DCT-based projection is a fast and accurate replacement for the orthogonal matrices computed via SVD/QR for the Adam-like optimizers (such as FRUGAL, FIRA, LDAdamW) in the context of low-rank compression of optimizer states. We reduce the running time and improve the memory usage as we store only one DCT matrix per GPU for the entire network, computed once at the beginning of the training. In addition, we store only $r$ (rank) integers, representing the indices of the most significant columns for each layer instead of storing one projection matrix.
@@
      DCT-AdamW (Section 2.4), which replaces the SVD-based low-rank projections and optionally adds quantized error feedback for the projection error, rotates the momentum buffers such that new low-rank gradients are correctly incorporated and allows changing the low-rank subspace every step;
@@ 2 Method
This section presents our dynamic column selection approach that works with any orthogonal matrix, briefly introduces the Discrete Cosine Transform (DCT) and the motivation of using it and finally introduce the two algorithms we propose: Trion as improvement to Dion to replace Power-Iteration and DCT-AdamW as an improvement to low-rank AdamW variants to replace SVD/QR.
@@ 2.1 Dynamic Column Selection
General View. Given an orthogonal matrix $Q\in\mathbb{R}^{n\times n}$, we need to compute a projection matrix $Q_{r}\in\mathbb{R}^{n\times r}$ by selecting $r$ columns from $Q$ to project the gradient $G\in\mathbb{R}^{n\times n}$ to an $r$-dimensional space $g=GQ_{r}\in\mathbb{R}^{n\times r}$. First, we compute the similarities matrix $S=GQ$ containing the scalar products between rows of $G$ and columns of $Q$. We rank the columns of $S$ based on their $\ell_{1}$- or $\ell_{2}$-norm and then pick the indices of the largest $r$ columns according to the chosen norm, which we use to index columns in $Q$ to obtain $Q_{r}$. The $i^{th}$ column in matrix $S$ contains the scalar products between each row of $G$ and $i^{th}$ column of $Q$, as detailed in Appendix B. We view these scalar products as similarities (or alignments) of rows in $G$ with columns of $Q$ and we want to choose the columns of $Q$ with largest similarity with rows of $G$. Using this approach, we ensure a dynamic mechanism to choose the most appropriate columns of $Q$ for the current gradient matrix $G$ to minimize the projection error (see Section 4). Each layer will have its own set of $r$ indices for the columns. The dynamic behavior comes from the $n$-choose-$r$ possible sets of indices to select from.
@@ 2.3 Trion: DCT-based improvement to Dion
In this section, we present how we can apply the DCT-based dynamic column selection approach to compute low-rank projection of momentum in the context of Dion (Ahn et al., 2025) to obtain a faster and more accurate optimizer, which we call Trion.
Dion uses Power-Iteration to compute a low-rank projection and orthogonalizes it using QR-decomposition, whose running time depends on the rank $r$. Instead, we propose replacing these techniques with a rank independent approach to compute the low-rank projection based on DCT matrix, which requires computing the similarity matrix $S_{t}$, followed by a top-$r$ ranking to determine the index of columns that best align with the momentum matrix, denoted by the set $i_{t}$. The matrix $S_{t}$ represents the DCT of the momentum matrix and it can be computed using the Makhoul’s algorithm or simply by only one matrix multiplication $S_{t}=B_{t}D_{C}$, where $B_{t}\in\mathbb{R}^{R\times C}$ is the momentum and $D_{C}\in\mathbb{R}^{C\times C}$ is the DCT matrix.
@@
We would like to emphasize that we input the low-rank momentum $b_{t}\in\mathbb{R}^{R\times r}$ to Newton-Schulz and not the original momentum buffer $B_{t}\in\mathbb{R}^{R\times C}$, which significantly reduces the FLOPs. Moreover, we can use the efficient triton kernels provided in the Muon implementation from the official Dion repository111github.com/microsoft/dion to speed up the computations even further for large ranks $r$.
Communication in Distributed Training. We base the Trion implementation on the published code for the Muon 222github.com/KellerJordan/Muon/blob/master/muon.py#L138 optimizer that leverages the ZeRO (Rajbhandari et al., 2020) approach in data-parallel settings. Specifically, the model update $O_{t}$ for a layer is computed only on one GPU, and the result is communicated to other GPUs using all-gather, drastically reducing the computation costs for large models. Since we replicate the DCT matrix on each GPU, we would like to emphasize that we communicate only low-rank terms $o_{t}$ from the source GPU to other devices and perform the step $O_{t}=o_{t}Q_{t}^{\top}$ locally on each device. This way, we do not communicate full size, orthogonal matrices $O_{t}$, but only low-rank versions $o_{t}$, thus reducing the overall iteration time.
@@
11:  $o_{t}=\textsc{NewtonSchulz}(b_{t})\in\mathbb{R}^{R\times r}$ $\triangleright$ Newton-Schulz applied on low-rank $b_{t}$, can use triton kernels
@@ 2.4 DCT-AdamW
We propose DCT-AdamW, a standalone low-rank version of AdamW with DCT-based projection that has the option to use quantized error feedback (EF) (Seide et al., 2014; Karimireddy et al., 2019; Alistarh et al., 2018) and ensures momentum buffers integrate gradients from the same lower dimensional subspaces in a similar way as in LDAdamW (Robert et al., 2025). In contrast to LDAdamW, which has to store two consecutive projection matrices per layer, we only have to store two sets of $r$ indices per layer. Prior work MicroAdam (Modoranu et al., 2024) quantized the error feedback down to 4-bits in the context of compressing the optimizer state using sparsity. In our setup, the lowest resolution we can quantize EF to is 8-bits without degrading the optimizer performance. or the space constraints reasons, we present the details of our DCT-AdamW optimizer in Appendix E.
@@ 3 Experiments
In this section we present our numerical results. Our main goal is to show that our DCT-based dynamic column selection approach at least recovers the performance of the original algorithms which we integrate it in. Specifically, we evaluate Trion against Dion, where we directly compare the DCT-based projection followed by Newton-Schulz iteration with Power-Iteration. In addition, we replace the SVD-based projection in FRUGAL and FIRA optimizers with our DCT approach. In the end, we compare LDAdamW with our standalone DCT-AdamW optimizer. In our pretraining experiments we compare training/validation perplexity and for fine-tuning we compare the evaluation accuracy, as well as memory usage and running time for both scenarios. In the remaining part of the paper we focus on presenting our pretraining (PT) results and the fine-tuning (FT) results are presented in Appendix H.
We train from scratch models from the Llama family with 350M, 800M and 1.3B parameters using Chinchilla-optimal token counts (20 tokens per parameter) from the C4 (Raffel et al., 2020) dataset and sequence length 512. All PT experiments are run in Distributed Data Parallel (DDP) settings on 8x H100 Nvidia GPUs using batch size 512 with local batch size 64 per GPU (unless otherwise specified explicitly).
PT with Trion. Our purpose is to show that our DCT-based dynamic column selection approach can successfully replace the QR-based Power-Iteration procedure in Dion and at least recovers the performance for the best hyper-parameters reported by the original work. In Table 1, we present our results, which are obtained using the optimal learning rate $\eta=0.01$ (as reported by the original Dion paper) and weight decay $\lambda=0.01$. We report the average values acros $3$ seeds for train/validation loss and perplexity, maximum allocated memory (read directly from the GPU) and the lowest running time across all runs. We experiment with ranks 128, 256 and 512 and the rank-to-dimension ratio is $r/d\in\{1/16,1/8,1/4,1/2\}$, equivalent to $6.25\%,12.5\%,25\%$ and $50\%$ of the full rank $d$, where $d$ is model’s embedding dimension.
@@
Table 1: Perplexity, Memory, and Running Time Comparison for Trion and Dion. $d$ stands for the embedding dimensionality of the model. The 1.3B model was trained with local batch size 32
Performance In Table 1, we show Trion consistently achieves lower training and validation loss, which also translates to lower perplexities. At the top row of Figure 3 we show the training loss curve of Trion is lower than Dion over time. This is an indication that our DCT-based dynamic column selection algorithm, followed by orthogonalization via Newton-Schulz can successfully replace the Power-Iteration procedure in Dion.
@@
Runtime. We would like to emphasize that running time of Trion does not depend on rank, as can be seen from the reported runtime across different ranks in Table 1: Trion achieves nearly constant runtime for each model size across all ranks, while the runtime of Dion clearly depends on the rank because of the QR-decomposition. This represents an advantage of Trion for much larger scales compared to Dion. At the bottom row of Figure 3 we present the wall clock time for the two optimizers for rank $256$ across all three models we tested on. Concretely, for a fixed running time budget, Trion consistently achieves lower training loss than Dion. Trion is faster than Dion by $2.5-4.5\%$ for rank $128$, $4.5-9\%$ for rank $256$ and $8-18\%$ for rank $512$. It is important to mention that the embedding size $d$ of the models we used in our work is too small to see a significant difference in the runtime between Makhoul’s algorithm and matmul when computing the column similarities. However, our benchmark presented in Appendix D shows the advantage of using Makhoul’s algorithm at scale compared to matmul.
@@
PT with LDAdamW. We train Llama-800M on 80B tokens (100 tokens per parameter) from C4 using LDAdam and DCT-AdamW described in Algorithm 2. In particular for our approach, we use EF quantized to 8-bits and some further optimizations from the ZeRO-redundancy optimizer (Rajbhandari et al., 2020), where one layer is updated on a single GPU and then it is broadcasted to the other GPUs. This way, we obtain lower memory usage by replacing redundant operations in the optimizer with communication, since the GPUs receiving the updated layer parameters do not allocate any state, at the cost of slightly increased running time due to more communication. In Figure 2 we present the training loss curves for full-rank AdamW (for reference), LDAdamW and DCT-AdamW and we are interested in directly comparing the last two. We observe that DCT-AdamW has lower training loss than LDAdamW, which also translates to lower perplexities in Table 2. Due to relatively high rank, the memory usage of LDAdamW is close to the memory usage of AdamW because it stores two projection matrices to be able to rotate the momentum buffers. In contrast, DCT-AdamW stores only two sets of $r$ indices instead of storing the actual projection matrices, which drastically reduces the memory usage, coupled with the ZeRO-redundancy trick. In terms of running time, DCT-AdamW is faster than LDAdamW by 10h 7m ($\approx 25.75\%)$ and slower than AdamW by 1h 55m ($\approx 5\%)$.
@@
PT with FRUGAL/FIRA. We replace the SVD projection with our DCT approach and test it on Llama-800M model using 16B tokens from the C4. For the space constraints reasons, we present the results in Appendix G.

### v3

@@ Abstract
Low-rank optimization has emerged as a promising direction in training large language models (LLMs) to improve running time and reduce the memory usage of adaptive optimizers by constraining learning to a lower-dimensional space. Prior work typically projects gradients of linear layers using approaches based on Singular Value Decomposition (SVD) or QR-decomposition. Applying these techniques individually to each layer in large models is computationally expensive and incurs additional memory costs due to storing the projection matrices. In this work, we propose a computationally efficient and conceptually simple, two-step procedure to approximate SVD/QR-based gradient projections into lower-dimensional spaces by using a predefined orthogonal matrix of the Discrete Cosine Transform (DCT). We dynamically select columns from the DCT matrix based on their alignment with the gradient of each layer. The effective projection matrices are obtained via a simple matmul with the DCT matrix in $O(n^{3})$ time, followed by a lightweight sorting step to identify the most relevant basis vectors. For large layers, DCT can be computed via Makhoul’s $N$-point algorithm based on Fast Fourier Transform (FFT) in $O(n^{2}\log(n))$ time. Due to the predefined nature of the orthogonal bases, they are computed once at the start of training. Our numerical experiments on both pre-training and fine-tuning tasks demonstrate the effectiveness of our dual strategy in approximating optimal low-rank projections, obtaining an approach with rank-independent running time that matches the performance of costly SVD/QR-based methods while achieving faster runtime and reduced memory usage by up to $25\%$ across different model sizes. Our code is available at ISTA-DASLab-Optimizers.
@@ 1 Introduction
Recently, the Muon (Jordan et al., 2024) optimizer sparked the community’s attention due to the faster convergence in pretraining settings, achieved by orthogonalizing the momentum matrix using Newton-Schulz, an iterative procedure difficult to parallelize for large-scale pretraining runs as the full-sized matrices must be materialized on GPU. The Dion optimizer (Ahn et al., 2025) aims to reduce this overhead by employing low-rank, orthogonal updates. However, it requires storing a projection matrix for each layer and uses QR-decomposition to orthogonalize the low-rank components, making its running time dependent on the rank.
@@
  • We motivate that DCT is a good candidate for our dynamic column selection approach due to the reduced time complexity to compute the alignments with the input matrix via the Makhoul’s $N$-point algorithm (Makhoul, 1980) to compute a fast DCT with $O(n^{2}\log(n))$ time complexity compared to $O(n^{3})$ incurred by a basic matrix multiplication. This can yield speedups of up to $8-50\times$ in computing the alignments for large layers (Appendix C).
@@
  • We show the DCT-based projection is a fast and accurate replacement for the orthogonal matrices computed via SVD/QR for the Muon- and Adam-like optimizers (such as Dion, FRUGAL, FIRA, LDAdamW) in the context of low-rank compression of optimizer states. We reduce the running time and improve the memory usage as we store only one DCT matrix per GPU for the entire network, computed once at the beginning of the training. In addition, we store only $r$ (rank) integers, representing the indices of the most significant columns for each layer instead of storing one low-rank projection matrix.
@@
      DCT-AdamW (Section 2.4), which replaces the SVD-based low-rank projections and optionally adds quantized error feedback for the projection error. Further, DCT-AdamW rotates the momentum buffers such that new low-rank gradients are correctly incorporated, and thus allows changing the low-rank subspace every step.
@@ 2 Method
This section presents our dynamic column selection approach that works with any orthogonal matrix, briefly introduces the Discrete Cosine Transform (DCT) and the motivation of using it and finally introduces the two algorithms we propose: Trion as an improvement to Dion to replace Power-Iteration and DCT-AdamW as an improvement to low-rank AdamW variants to replace SVD/QR.
@@ 2.1 Dynamic Column Selection
General View. Given an orthogonal matrix $Q\in\mathbb{R}^{n\times n}$, we want to compute a projection matrix $Q_{r}\in\mathbb{R}^{n\times r}$ by selecting $r$ columns from $Q$ to project the gradient $G\in\mathbb{R}^{n\times n}$ to an $r$-dimensional space $g=GQ_{r}\in\mathbb{R}^{n\times r}$. First, we compute the similarities matrix $S=GQ$ containing the scalar products between rows of $G$ and columns of $Q$. We rank the columns of $S$ based on their $\ell_{1}$- or $\ell_{2}$-norm and then pick the indices of the largest $r$ columns according to the chosen norm, which we use to index columns in $Q$ to obtain $Q_{r}$. The $i^{th}$ column in matrix $S$ contains the scalar products between each row of $G$ and $i^{th}$ column of $Q$, as detailed in Appendix B. We view these scalar products as similarities (or alignments) of rows in $G$ with columns of $Q$ and we want to choose the columns of $Q$ with largest similarity with rows of $G$. Using this approach, we ensure a dynamic mechanism to choose the most appropriate columns of $Q$ for the current gradient matrix $G$ to minimize the projection error (see Section 4). Each layer will have its own set of $r$ indices for the columns. The dynamic behavior comes from the $n$-choose-$r$ possible sets of indices to select from.
@@ 2.3 Trion: DCT-based improvement to Dion
In this section, we present how we can apply the DCT-based dynamic column selection approach to compute low-rank projection of momentum in the context of Dion (Ahn et al., 2025) to obtain a faster and more accurate optimizer, which we call Trion, presented in Algorithm 1.
Dion uses Power-Iteration to compute a low-rank projection and orthogonalizes it using QR-decomposition, whose running time depends on the rank $r$. Instead, we propose replacing these techniques with a rank independent approach to compute the low-rank projection based on DCT matrix, which requires computing the similarity matrix $S_{t}$, followed by a top-$r$ ranking to determine the indices of columns that best align with the momentum matrix, denoted by the set $i_{t}$. The matrix $S_{t}$ represents the DCT of the momentum matrix and it can be computed using the Makhoul’s algorithm or simply by only one matrix multiplication $S_{t}=B_{t}D_{C}$, where $B_{t}\in\mathbb{R}^{R\times C}$ is the momentum and $D_{C}\in\mathbb{R}^{C\times C}$ is the DCT matrix.
@@
We would like to emphasize that we input the low-rank momentum $b_{t}\in\mathbb{R}^{R\times r}$ to Newton-Schulz and not the original momentum buffer $B_{t}\in\mathbb{R}^{R\times C}$, which significantly reduces the computational overhead. Moreover, we can use the efficient triton kernels provided in the Muon implementation from the official Dion repository111github.com/microsoft/dion to speed up the computations even further for large ranks $r$, as Newton-Schulz will operate with $r\times r$ matrices.
Communication in Distributed Training. We develop the Trion optimizer on top of the published code for the Muon  [^2] Keller Jordan’s Muon implementation optimizer that leverages the ZeRO (Rajbhandari et al., 2020) approach in Distributed Data Parallel (DDP) settings. Specifically, the model update $O_{t}$ for a layer is computed only on one GPU, and the result is communicated to other GPUs using all-gather, drastically reducing the computation costs for large models (under the assumption that communication is cheaper than computation itself). Since we replicate the DCT matrix on each GPU, we would like to emphasize that we communicate only low-rank terms $o_{t}$ from the source GPU (the one computing the update) to other devices and perform the step $O_{t}=o_{t}Q_{t}^{\top}$ locally on each device. This way, we do not communicate full size, orthogonal matrices $O_{t}$, but only low-rank versions $o_{t}$, thus reducing the overall iteration time.
@@
11:  $o_{t}=\textsc{NewtonSchulz}(b_{t})\in\mathbb{R}^{R\times r}$ $\triangleright$ Newton-Schulz applied on low-rank $b_{t}$
@@ 2.4 DCT-AdamW
We propose DCT-AdamW, a standalone low-rank version of AdamW with DCT-based projection that has the option to use quantized error feedback (EF) (Seide et al., 2014; Karimireddy et al., 2019; Alistarh et al., 2018) and ensures momentum buffers integrate gradients from the same lower dimensional subspaces in a similar way as in LDAdamW (Robert et al., 2025). In contrast to LDAdamW, which has to store two consecutive projection matrices per layer, we only have to store two sets of $r$ indices per layer. Prior work MicroAdam (Modoranu et al., 2024) quantized the error feedback down to 4-bits in the context of compressing the optimizer state using sparsity. In our setup, the lowest resolution we can quantize EF to is 8-bits without degrading the optimizer performance. For space constraints reasons, we present the pseudocode of our DCT-AdamW optimizer in Appendix E.
@@ 3 Experiments
In this section we present our numerical results. Our main goal is to show that our DCT-based dynamic column selection approach at least recovers the performance of the original algorithms which we integrate it in. Specifically, we evaluate Trion against Dion, where we directly compare the DCT-based projection followed by Newton-Schulz iteration with QR-based Power-Iteration. In addition, we replace the SVD-based projection in FRUGAL and FIRA optimizers with our DCT approach. In the end, we compare LDAdamW with our standalone DCT-AdamW optimizer. In our pretraining experiments, we compare training/validation perplexity and for fine-tuning we compare the evaluation accuracy, as well as memory usage and running time for both scenarios. In the remaining part of the paper, we focus on presenting our pretraining (PT) results and the fine-tuning (FT) results are presented in Appendix H.
We train from scratch models from the Llama family with 350M, 800M and 1.3B parameters using Chinchilla-optimal token counts (20 tokens per parameter) from the C4 (Raffel et al., 2020) dataset and sequence length 512. All PT experiments are run in Distributed Data Parallel (DDP) settings on 8x H100 Nvidia GPUs using global batch size 512 with local batch size 64 per GPU (unless otherwise specified explicitly).
PT with Trion. Our purpose is to show that our DCT-based dynamic column selection approach can successfully replace the QR-based Power-Iteration procedure in Dion and at least recovers the performance for the best hyper-parameters reported by the original work. In Table 1, we present our results, which are obtained using the optimal learning rate $\eta=0.01$ (as reported by the original Dion paper) and weight decay $\lambda=0.01$. We report the average values across $3$ seeds for train/validation loss and perplexity, maximum allocated memory (read directly from the GPU) and the lowest running time across all runs. We experiment with ranks $128,256$ and $512$, the rank-to-dimension ratio is $r/d\in\{1/16,1/8,1/4,1/2\}$, equivalent to $6.25\%,12.5\%,25\%$ and $50\%$ of the full rank $d$, where $d$ is model’s embedding dimension.
@@
Table 1: Perplexity, Memory, and Running Time Comparison for Trion and Dion. $d$ stands for the embedding dimensionality of the model. The 1.3B model was trained with local batch size 32.
Performance In Table 1, we show Trion consistently achieves lower training and validation loss, which also translates to lower perplexities. At the top row of Figure 3 we show the training loss curve of Trion is lower than Dion across iterations. This is an indication that our DCT-based dynamic column selection algorithm, followed by orthogonalization via Newton-Schulz can successfully replace the Power-Iteration procedure in Dion.
@@
Runtime. We would like to emphasize that the running time of Trion does not depend on rank, as can be seen from the reported runtime across different ranks in Table 1: Trion achieves nearly constant runtime for each model size across all ranks, while the runtime of Dion clearly depends on the rank because of the QR-decomposition. This represents an advantage of Trion for much larger scales compared to Dion. At the bottom row of Figure 3 we present the wall clock time for the two optimizers for rank $256$ across all three models we tested on. Concretely, for a fixed running time budget, Trion consistently achieves lower training loss than Dion. Trion is faster than Dion by $2.5-4.5\%$ for rank $128$, $4.5-9\%$ for rank $256$ and $8-18\%$ for rank $512$ (the overhead of Dion increases with rank). It is important to mention that the embedding size $d$ of the models we used in our work is too small to see a significant difference in the runtime between Makhoul’s algorithm and matmul when computing the column similarities. However, our benchmark presented in Appendix D shows the advantage of using Makhoul’s algorithm at scale compared to matmul.
@@
PT with LDAdamW. We train Llama-800M on 80B tokens (100 tokens per parameter) from C4 using LDAdam and DCT-AdamW described in Algorithm 2. In particular for our approach, we use EF quantized to 8-bits and some further optimizations from the ZeRO-redundancy optimizer (Rajbhandari et al., 2020), where one layer is updated on a single GPU and then it is broadcasted to the other GPUs. This way, we obtain lower memory usage by replacing redundant operations in the optimizer with communication, since the GPUs receiving the updated layer parameters do not allocate any state. In Figure 2 we present the training loss curves for full-rank AdamW (for reference), LDAdamW and DCT-AdamW and we are interested in directly comparing the last two. We observe that DCT-AdamW has lower training loss than LDAdamW, which also translates to lower perplexities in Table 2. Due to relatively high rank, the memory usage of LDAdamW is close to the memory usage of AdamW because it stores two projection matrices to be able to rotate the momentum buffers. In contrast, DCT-AdamW stores only two sets of $r$ indices instead of storing the actual projection matrices, which drastically reduces the memory usage, coupled with the ZeRO-redundancy trick. In terms of running time, DCT-AdamW is faster than LDAdamW by 10h 7m ($\approx 25.75\%)$ and slower than AdamW by 1h 55m ($\approx 5\%)$.
@@
PT with FRUGAL/FIRA. We replace the SVD projection with our DCT approach and test it on Llama-800M model using 16B tokens from the C4. For space constraints reasons, we present the results in Appendix G.